# New Perspectives in Advancing Graph Machine Learning

### v1

@@ E.2.4 GCN, GIN, GAT
E.2.5 EGNN and TFN
For the node-level equivariant regression task, we tune the number of layers separately for each task by starting from one and increasing until performance degrades rather than improves. The best-performing configurations use two layers for EGNN and four layers for TFN on the graph-level task; and seven and four layers, respectively, on the node-level task. Prediction heads are implemented as two-layer MLPs (which operate on both sum and max pooled hidden node features for graph-level predictions).
@@ E.2.5 EGNN and TFN
E.3 Details on the node-level vector target
@@
and rescale $\mathbf{g}$ so that its entries have magnitudes in $[0,a]$, for some fixed $0<a<1$, by setting $\widetilde{\mathbf{g}}=a\frac{\mathbf{g}}{{\|\mathbf{g}\|_{\infty}}}$. Finally, we define the magnitude of $\mathbf{h}$ by $\|\mathbf{h}(v)\|_{2}=1+\widetilde{\mathbf{g}}(v).$ We note that by construction, we will have $0<1-a\leq\|\mathbf{h}(v)\|_{2}\leq 1+a$ for all $v$.
@@ E.4 Model runtimes
The run times for each model are for the diameter prediction task and the node-level vector prediction task are displayed in Tables 3 and 4 respectively. We note that all models were trained using one NVIDIA L40S 48 GB GPU, with one exception. Due to its high parameter count and larger memory requirements, we trained TFN using distributed data parallel (DDP) with four such GPUs. Hence we conjecture that, if the model fit on one GPU, its runtimes would roughly quadruple.
@@
Table 3: Epoch number of best validation loss, average training epoch runtime, and validation set inference time on the diameter prediction task. TFN is marked with **** since it was trained on multiple GPUs (while all other models used a single GPU), as explained at the beginning of section E.4.
Model Best epoch Avg. train time (s/epoch) Avg. inference time (s)
ESc-GNN (Ours) $461\pm\;\;56$ $0952\pm 0.0008 0.0$$0168\pm 0.0001$
ESc-GNN (non-equivariant) (Ours)   $96\pm\;\;93$ $0792\pm 0.0039 0.0$$0072\pm 0.0034$
LEGS $356\pm 155$ $0946\pm 0.0014 0.0$$0164\pm 0.0002$
GCN $335\pm 132$ $0568\pm 0.0009 0.0$$0038\pm 0.0000$
GAT   $62\pm\;\;31$ $0662\pm 0.0013 0.0$$0048\pm 0.0000$
GIN $130\pm 111$ $0576\pm 0.0004 0.0$$0027\pm 0.0000$
EGNN (2-layer) $338\pm 178$ $0856\pm 0.0009 0.0$$0074\pm 0.0001$
TFN (4-layer)**** $453\pm\;\;72$ $2.6965\pm 0.0044$ $0.0880\pm 0.0007$
@@
Table 4: Epoch number of best validation loss, average training epoch runtime, and validation set inference time on the node-level vector prediction task. TFN is marked with **** since it was trained on multiple GPUs (while all other models used a single GPU), as explained at the beginning of section E.4.
Model Best epoch Avg. train time (s/epoch) Avg. inference time (s)
ESc-GNN (Ours) $488\pm\;\;\;\;4$ $0945\pm 0.0003 0.0$$0164\pm 0.0000$
ESc-GNN (non-equivariant) (Ours) $355\pm 184$ $0793\pm 0.0034 0.0$$0101\pm 0.0042$
LEGS $402\pm 128$ $0980\pm 0.0016 0.0$$0163\pm 0.0001$
GCN $356\pm 176$ $0583\pm 0.0007 0.0$$0035\pm 0.0000$
GAT $356\pm 179$ $0624\pm 0.0004 0.0$$0042\pm 0.0000$
GIN $330\pm 171$ $0562\pm 0.0004 0.0$$0021\pm 0.0000$
EGNN (7-layer) $325\pm\;\;77$ $1552\pm 0.0015 0.0$$0183\pm 0.0001$
TFN (4-layer)**** $364\pm 123$ $2.6912\pm 0.0042$ $0.0884\pm 0.0017$

### v2

@@ E.3.4 GCN, GIN, GAT
E.3.5 EGNN and TFN
For the EGNN and TFN models, we adapted code from the “Geometric GNN Dojo” (Joshi et al., 2023) under the MIT license, from their Github repository (https://github.com/chaitjo/geometric-gnn-dojo). For these models, on the node-level equivariant regression task, we tune the number of layers separately for each task by starting from one and increasing until performance degrades rather than improves. The best-performing configurations use two layers for EGNN and four layers for TFN on the graph-level task; and seven and four layers, respectively, on the node-level task. Prediction heads are implemented as two-layer MLPs (which operate on both sum and max pooled hidden node features for graph-level predictions).
@@ E.3.5 EGNN and TFN
E.4 Details on the node-level vector target
@@
and rescale $\mathbf{g}$ so that its entries have magnitudes in $[0,a]$, for some fixed $0<a<1$, by setting $\widetilde{\mathbf{g}}=a\frac{\mathbf{g}}{{\|\mathbf{g}\|_{\infty}}}$. Finally, we define the magnitude of $\mathbf{h}$ by $\|\mathbf{h}(v)\|_{2}=1+\widetilde{\mathbf{g}}(v).$ We note that by construction, we have $0<1-a\leq\|\mathbf{h}(v)\|_{2}\leq 1+a$ for all $v$.
@@ E.5 Model runtimes
The run times for each model are for the diameter prediction task and the node-level vector prediction task are displayed in Tables 5 and 6 respectively. We note that all models were trained using one NVIDIA L40S 48 GB GPU, with one exception. Due to its high parameter count and larger memory requirements, we trained TFN using distributed data parallel (DDP) with four such GPUs. Hence we conjecture that, if the model fit on one GPU, its runtimes would roughly quadruple.
@@
Table 5: Epoch number of best validation loss and average training epoch runtime on the diameter prediction task. TFN is marked with [4 GPUs] since it was trained on multiple GPUs (while all other models used a single GPU), as explained at the beginning of section E.5.
Model Best epoch Sec. per epoch
VDW-GNN (Ours) $461\pm 56$ $0.0952\pm 0.0008$
VDW-GNN (non-equivariant) (Ours) $96\pm 93$ $0.0792\pm 0.0039$
LEGS $356\pm 155$ $0.0946\pm 0.0014$
GCN $335\pm 132$ $0.0568\pm 0.0009$
GAT $62\pm 31$ $0.0662\pm 0.0013$
GIN $130\pm 111$ $0.0576\pm 0.0004$
EGNN (2-layer) $338\pm 178$ $0.0856\pm 0.0009$
TFN (4-layer) [4 GPUs] $453\pm 72$ $2.6965\pm 0.0044$
@@
Table 6: Epoch number of best validation loss and average training epoch runtime on the node-level vector prediction task. TFN is marked with [4 GPUs] since it was trained on multiple GPUs (while all other models used a single GPU), as explained at the beginning of section E.5.
Model Best epoch Sec. per epoch
VDW-GNN (Ours) $488\pm 4$ $0.0945\pm 0.0003$
VDW-GNN (non-equivariant) (Ours) $355\pm 184$ $0.0793\pm 0.0034$
LEGS $402\pm 128$ $0.0980\pm 0.0016$
GCN $356\pm 176$ $0.0583\pm 0.0007$
GAT $356\pm 179$ $0.0624\pm 0.0004$
GIN $330\pm 171$ $0.0562\pm 0.0004$
EGNN (7-layer) $325\pm 77$ $0.1552\pm 0.0015$
TFN (4-layer) [4 GPUs] $364\pm 123$ $2.6912\pm 0.0042$
Appendix F Experimental details - Wind field reconstruction (Section 4.2)
Lifting 2D wind vectors to 3D. Given horizontal surface wind components $\mathbf{w}_{\text{2D}}=(u,v)$ (eastward, northward), for a point on the unit sphere with latitude $\phi$ and longitude $\lambda$, we compute the 3D wind vector by $\mathbf{w}_{\text{3D}}=u\,{\mathbf{e}}_{\text{east}}+v\,{\mathbf{e}}_{\text{north}}$ where ${\mathbf{e}}_{\text{east}}=[-\sin\lambda,\cos\lambda,0]^{\top},$ and ${\mathbf{e}}_{\text{north}}=[-\sin\phi\cos\lambda,-\sin\phi\sin\lambda,\cos\phi]^{\top}$ are basis vectors for the local tangent space.
Architecture. Here, our VDW-GNN is a lightweight vector-only graph module that maps per-node vectors to updated vectors in each block by (1) applying vector diffusion wavelets; (2) concatenating the resulting coefficients with the ordered top-$k$ incoming neighbor-vector and edge-weight features, and (3) feeding into a shared MLP to predict a new vector for each node. In the wind configuration, a single block is used with a residual update on the original vector feature. We use a custom set of vector diffusion wavelets $\widetilde{\mathcal{W}}_{J}=\{(\mathbf{I}-\mathbf{Q}),(\mathbf{Q}-\mathbf{Q}^{2}),(\mathbf{Q}^{2}-\mathbf{Q}^{3}),\mathbf{Q}^{3}\}$ (where the last is the low-pass wavelet).
Training procedure. We enforced the following early stopping procedure for all models: after a burn-in of at least 100 epochs, if validation MSE (calculated every epoch) does not improve for 100 epochs, halve the learning rate and continue training from the previous best model weights by validation MSE, and then cease training if validation MSE fails to improve for another 100 consecutive epoch period.
Hyperparameter settings. Tuning the number of layers for GIN, GAT, GCN, and EGNN, we found that a single layer was optimal for these models; and two for TFN. We found a node embedding dimension of 128 worked best for EGNN, and 16 for TFN. We used a maximum irreducible representation order of $\ell=2$ for TFN. All models use a final three-layer prediction MLP head with hidden widths (128, 128) and output dimension 3, SiLU activations, and no batch normalization or dropout. For DD-TNN, we maintained hyperparameter settings as set by Battiloro et al. (2024) in their codebase. We trained all models with PyTorch’s AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.999)$, weight decay of $10^{-6}$, and an initial learning rate of 0.005.
Appendix G Experimental details - Multi-channel neural recordings (Section 4.3)
Figure 5: Models’ distributions of mean (of mean cross-validated) classification accuracies across $4$ experimental days. Brackets annotation the $p$-value results of two-sided, paired Wilcoxon tests (ns: not significant, ****: $p<0.0001$).
Data processing. We use the data set version available on MARBLE’s data repository (Gosztolai, 2023), which is a processed subset of the original set of experiments in Kaufman et al. (2016), and in which the raw trial recordings have been converted into rates using a Gaussian kernel (s.d. 100 ms), then subsampled to 20-millisecond intervals. [^6] In MARBLE’s version of this reduced data set, some days have seven conditions, while others have eight; however, MARBLE used seven in their analysis, and we follow suit. Following MARBLE’s data processing protocol, for all models, we truncate the time series to after the ‘go’ cues, leaving 35 timepoints per trial, and then smooth with a Savitzky-Golay filter (order 2, window length 9). However, unlike MARBLE, we do not then apply PCA to reduce the data dimension below 24 (which we found discards useful signal, as a severe linear reduction). For our model, we follow MARBLE’s lead, and reconstruct the problem as learning on a graph with vector-valued node features (note that CEBRA and LFADS do not convert the neural data to vector-valued features). Here, we follow the convention of assigning velocity vectors derived as $\mathbf{v}_{t}=\mathbf{x}_{t+1}-\mathbf{x}_{t}$ to timepoint $t$ and dropping the last timepoint, leaving a final 34 nodes per trial.
Baselines. MARBLE is a graph-based deep learning method that learns embeddings of neural states conceived as local flow fields on a manifold, using gradient filters and unsupervised contrastive loss. CEBRA combines nonlinear independent components analysis (ICA) and self-supervised contrastive learning to produce embeddings that jointly model neural activity data with behavioral auxiliary variables and temporal structure. LFADS is an unsupervised method built from sequential autoencoders, which instead of contrastive loss, uses a reconstruction training objective and KL divergence regularization. Additionally, we note that while MARBLE (and our model) compute and use vector-valued neural ‘velocity’ features, CEBRA and LFADS do not use vector-valued features, but instead ingest scalar-valued neural recording data.
Architecture. Our model for this task consists of three layers: (1) an equivariant, vector-valued geometric scattering transform of timepoints’/nodes’ neural velocity vector features (zeroth- and first-order transforms only, using a custom set of vector diffusion wavelets $\widetilde{\mathcal{W}}_{J}=\{(\mathbf{I}-\mathbf{Q}),(\mathbf{Q}-\mathbf{Q}^{2}),(\mathbf{Q}^{2}-\mathbf{Q}^{3}),\mathbf{Q}^{3}\}$ (where the last is the low-pass wavelet); (2) a four-layer MLP that projects the concatenated multi-order scattering coefficients into the embedding dimension (with structure input $\rightarrow 256,256,256\rightarrow$ embedding, ReLU activations, and slight dropout $p=0.05$); (3) a custom supervised contrastive loss function that treats nodes from the same behavioral condition as positives and nodes from different conditions as negatives, and focuses learning on informative “hard” negatives. Concretely, for each anchor embedding, we sample a fixed number of positives per anchor and then select a top‑$k$ subset of negatives with highest similarity (plus optional random negatives) to form a compact but challenging contrastive batch from which to compute loss as a temperature-scaled log-softmax over these similarities that encourages tight within-condition clusters and separation across conditions.
Inductive evaluation. A key contribution of our experimental design is to test models on inductive (as opposed to transductive) representation learning. This evaluation protocol is more stringent than the original MARBLE setup. Whereas MARBLE trained and evaluated in a transductive manner by randomly holding out points while still permitting the test nodes to appear in the graph during embedding computation, we withhold entire trials within each day. Thus models are trained only on training trials; test trials are never used in training nor incorporated into the training graph. This trial-wise split prevents subtle data leakage via neighborhood overlap, and directly probes generalization to unseen trial trajectories.
Across all methods, we evaluate representations via an SVM probe trained on training-trial embeddings and applied to held-out trial embeddings, ensuring that the decoding stage reflects the generalization properties of the learned representations rather than the capacity of an end-to-end classifier. In our inductive protocol, entire trials are withheld from training and never included when building the training graph for our model and for MARBLE (both of which are graph-based). This creates the challenge that test nodes do not exist in the training graph at evaluation time. Rather than reconstructing an augmented graph containing test nodes (which would revert to a semi-transductive setting), we estimate test-trial embeddings using a neighbor-based distillation procedure. That is, each test node’s embedding is computed as a Gaussian kernel–weighted average of the embeddings of its nearest neighbors in the training set, where neighbors are selected according to distance in neural activity space. This yields an inductive mapping from unseen nodes to the learned representation space without retraining or graph augmentation. In contrast, for CEBRA and LFADS do not use graphs or vector features. These models learn embeddings directly from trial time series, and test-trial embeddings are obtained by a forward pass through their trained encoder modules on the held-out test trials.
Training procedure. We employed the same learning rate adjustment and early stopping procedure as used in the wind experiments (Appendix F). However, the stopping rule is here governed by the classification accuracy achieved by the SVM (which is fit to the training set trials before each post-epoch evaluation) failing to improve for 32 epochs, after a burn-in of at least 32 epochs. For MARBLE, we preserved the learning rate reduction factor in their codebase as 0.1 (instead of 0.5, as we used). Note that CEBRA trains with its own fit() or partial_fit() methods, and we used partial_fit() to enforce early stopping by checking SVM validation accuracy every 100 iterations, as one epoch.
Hyperparameters. For the CkNN graph used by our model and MARBLE, we used $k=30$ neighbors and $\delta=1.0$ (note that this creates a sparser graph than MARBLE’s original $\delta=1.4$). The support vector machine (SVM) classifier used to evaluate embeddings used a radial basis function kernel, with scale parameter $\gamma=(p\sigma)^{-1}$ (where $p$ is the number of features and $\sigma$ is the overall variance of the training data), and a regularization constant of $1$.
For our supervised contrastive loss function, we used a log-softmax temperature of 0.05, and randomly sampled 128 ‘anchor’ nodes in each epoch, and further sampled one positive-sample node, eight top-$k$ neighboring negative sample nodes, and eight random (uniformly sampled) negative sample nodes per anchor.
LFADS has a large number of hyperparameters; our tuning focused on the ‘initial condition’ and ‘controller’ encoders widths (both ultimately set to 256). Additionally, because we processed the data into smoothed, continuous-valued signals (not counts), we set its reconstruction parameter to ‘‘gaussian" instead of ‘‘poisson".
We trained all models with PyTorch’s AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.999)$, weight decay of $10^{-5}$, and an initial learning rate of 0.005.